Calculate delta H degree for the process


2Al(s)+Fe2O3(s)=Al2O3(s)+2Fe(s)

Given that delta Hf degree of Fe2O3=-813.0KJ/mole and delta Hf degree of Al2O3 is -1655.0KJ/mole


1
Expert's answer
2021-06-12T10:42:09-0400

delta H degree for the process = Hf(products) - Hf(reagents) = Hf(Al2O3) - Hf(Fe2O3) = - 1655 - (-813) = - 842 KJ/mole